Wildhorse, Inc., has bonds outstanding that will mature in 8 years. The bonds have a face value of $1,000. These bonds pay interest semiannually and have a coupon rate of 4.6 percent. If the bonds are currently selling at $909.92, what is the yield to maturity that an investor who buys them today can expect to earn? What is the effective annual yield?
1
Expert's answer
2019-11-12T17:05:37-0500

Data;

Par value/ Face value= $1000

Coupon rate per year=4.6%

Coupon rte per semi annual period= "\frac{4.6\%} {2}" = 2.30%

Coupon payment= 0.023*1000=$23

Current price/ Present value=$909.92

Time= 8 years

Semi annual periods= 8*2=16


Calculation yield To Maturity

"YTM=\cfrac{C+\cfrac {(FV-PV)}{t}} {\cfrac{(FV+PV)} {2}}"


C=Coupon payment

FV= Face value

PV= Present value/ Price

t= No of years takes to maturity

"YTM=\cfrac{23+\cfrac {(1000-909.92)}{16}} {\cfrac{(1000+909.92)} {2}}"


"YTM=0.02998"

"YTM=2.998\%"

"YTM=0.02998"

"YTM=2.998\% * 2=5.996\%"

"YTM=\boxed{5.996\%}"


Calculation for Effective Annual Yield


"ENY=\lparen {1+\cfrac{r} {n}}\rparen ^{n} -1"


r= Interest rate

n=Number of payments per year


"ENY=\lparen {1+\cfrac{0.023} {2}}\rparen ^{2} -1"


"ENY=0.0231"


"ENY=2.31\%"

"ENY=2.31\% *2= 4.63\%"

"ENY=\boxed{4.63\%}"